OpenStudy (anonymous):

Solve for x showing steps: 27e^[x^(3)+4] = e^7

OpenStudy (helder_edwin):

take natural logarithms on both sides

OpenStudy (anonymous):

ln27e^[x^(3)+4] = lne^7 I understand that the exponents are dropped down, but what happens to the 27 in front of the e?

OpenStudy (helder_edwin):

\[\large \ln(AB)=\ln A+\ln B\] u have \[\large 27e^{x^3+4}=e^7 \] from this \[\large \ln(27e^{x^3+4})=\ln e^7 \]

OpenStudy (anonymous):

divide by \(27\) first

OpenStudy (anonymous):

\[x^3+4=\ln(\frac{e^7}{27})\]

OpenStudy (isaiah.feynman):

Alternatively, divide both sides by 27. you get \[e^{x^{3}+4}=\frac{ e^{7} }{ 27}\]Take natural logs of both sides\[x^{3}+4=\ln \left( \frac{ e^{7} }{ 27 } \right)\]Using logarithm rules on the right side we have\[x^{3}+4=\ln e^{7}-\ln27\]The right side simplifies to \[x^{3}+4=7-\ln27\] You can now solve for x from here.

OpenStudy (anonymous):

then you can write \[\ln(e^7)-\ln(27)=7-\ln(27)\]

OpenStudy (anonymous):

and don't forget \(3-\ln(27)\) is just some number, so you can solve via taking the cubed root

OpenStudy (anonymous):

Thank you guys so much!